OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!! Which of the following would result from multiplying each side of the polygon by a ratio of 1/2? The final area is 12km2 The final perimeter is 24km The final area is 6km2 The final perimeter is 20km

OpenStudy (anonymous):

please help me

OpenStudy (vishweshshrimali5):

yes

OpenStudy (vishweshshrimali5):

the final area will reduce

OpenStudy (vishweshshrimali5):

I cannot understan what u want ?

OpenStudy (anonymous):

The perimeter will be one-half that of the original polygon. EX: If the length of the sides of the polygon are w,x,y,z the original perimeter would be \[P = w+x+y+z\] And if you decrease the length by half of all the sides you would get \[P = {1\over 2}w+{1\over2}x+{1\over2}y+{1\over2}z\] \[P = {1\over2}\left(w+x+y+z\right)\] As for the area, it should decrease by a factor of 1/4. I'll use the example of a square with sides of length x. \[A = x^{2}\] The new area would be \[A = \left({1\over2}x*{1\over2}x\right)\] \[A = {1\over4}x^{2}\]

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

Depends on the area and perimeter of the original polygon.......which you haven't given us

OpenStudy (anonymous):

i only have the info i put up

OpenStudy (anonymous):